OpenStudy (anonymous):

When using substitution to solve the system below, which expression can be substituted into the second equation for y. 4x+y=3 and 2x+3y=4 a. 4-2x/3 b. 3-4x c. 4-2x

OpenStudy (kropot72):

Can you rearrange the first equation to find an expression for y ?

OpenStudy (anonymous):

idk how.

OpenStudy (kropot72):

4x + y = 3 What do you get when you subtract 4x from both sides of this equation?

OpenStudy (anonymous):

0+y=-1

OpenStudy (kropot72):

4x + y = 3 Subtractin 4x from both sides gives 4x - 4x + y = 3 - 4x Can you find an expression for y now?

OpenStudy (anonymous):

3 ?

OpenStudy (kropot72):

4x - 4x + y = 0 + y = y = 3 - 4x y = 3 - 4x